Assume a perfectly competitive firm's short-run cost is TC = 100 + 160Q + 3Q2. If the market price is ETB 196, How much should it produce and what should it do for long run production? 


1
Expert's answer
2021-08-29T16:50:34-0400
"TC=100+160Q+3Q^2"

"MC=160+6Q"

"160+6Q=196"

"6Q=36"

"Q=6"